# Global minimality of the Hopf map in the Faddeev–Skyrme model with large coupling constant

### v1

@@ 4. Relaxed form of the perturbed Faddeev–Skyrme energy
The next lemma gives us a precise expansion for the energy $\mathcal{E}_{\rho}$ around elements of that first eigenspace.
@@
$\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0,1})$ $=\rho^{-2}\bigg{(}\int_{\mathbb{S}^{3}}|d\varphi|^{2}-2\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi\bigg{)}-\frac{1}{2}\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi+\frac{1}{128\pi^{2}}\bigg{(}\int_{\mathbb{S}^{3}}\langle\alpha^{+}_{0,1},d\varphi\rangle\bigg{)}^{2}$
$\quad+\bigg{(}2-\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\langle\alpha^{+}_{0,1},d\varphi\rangle\bigg{)}\mathcal{G}_{1}(\alpha^{+}_{0,1},d\varphi)+\big{(}2+\rho^{-2}\big{)}\mathcal{O}\left(\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{3}\right),$ (4.12)
where we set
@@
$\mathcal{E}_{\rho}(\alpha)=2Q(\alpha)^{-1/2}\mathcal{I}_{1}(\alpha)+\rho^{-2}Q(\alpha)^{-1}\mathcal{I}_{2}(\alpha)\,.$ (4.15)
We first expand the three quantities appearing in (4.15) separately, obtaining:
$\begin{split}{\rm(i)}&\quad\mathcal{I}_{1}(\alpha)=\mathcal{I}_{1}(\alpha^{+}_{0,1})+\mathcal{I}_{1}^{\prime}(\alpha^{+}_{0,1})[d\varphi]+\mathcal{G}_{1}(\alpha^{+}_{0,1},d\varphi)\,,\\[2.0pt] {\rm(ii)}&\quad\mathcal{I}_{2}(\alpha)=\mathcal{I}_{2}(\alpha^{+}_{0,1})+\mathcal{I}_{2}^{\prime}(\alpha^{+}_{0,1})[d\varphi]+\int_{\mathbb{S}^{3}}|d\varphi|^{2}\,,\\[2.0pt] {\rm(iii)}&\quad Q(\alpha)^{-q/2}=\bigg{(}1+Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]+\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi\bigg{)}^{-q/2}\\[2.0pt] &\qquad\qquad\quad=1-\frac{q}{2}Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]-\frac{q}{32\pi^{2}}\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi+\frac{q(q+2)}{8}\big{(}Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]\big{)}^{2}\\[2.0pt] &\qquad\qquad\qquad\qquad+\mathcal{O}(\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{3})\,,\end{split}$ (4.16)
@@
$\mathcal{I}^{\prime}_{1}(\alpha^{+}_{0,1})[d\varphi]:=\frac{d}{dt}\Big{|}_{t=0}\mathcal{I}_{1}(\alpha^{+}_{0,1}+td\varphi)\,,$
and similarly for $\mathcal{I}_{2}^{\prime}(\alpha^{+}_{0,1})[d\varphi]$ and $Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]$. Note that (4.16)(i) follows simply by the definition of $\mathcal{G}_{1}$ in (4.13), and (4.16)(ii) by a simple expansion of the quadratic energy $\mathcal{I}_{2}$. For the first line in the third expansion we have used (4.10), the integral formula (4.1), together with $Q(\alpha^{+}_{0,1})=1$, cf. the definition of $E^{+}_{0,1}$ in (4.5). Finally, for the last line in (4.16)(iii) we used a Taylor expansion for $t\mapsto(1+t)^{-q/2}$ around 0, together with the estimate
$\Big{|}\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi\Big{|}=\mathcal{O}(\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{2})\,,$
coming from Lemma 2.1(ii).
To proceed further we note that, by the definition of $\beta_{0}$ in (4.10) and Lemma 2.1(i) we have
@@
$\mathcal{I}_{q}^{\prime}(\alpha^{+}_{0,1})=\frac{q}{2}\mathcal{I}_{q}(\alpha^{+}_{0,1})Q^{\prime}(\alpha^{+}_{0,1})\,;$ (4.18)
this last equation, together with $Q(\alpha^{+}_{0,1})=1$, expresses the fact that $\alpha^{+}_{0,1}$ is a critical point of the functionals $\alpha\mapsto Q^{-q/2}(\alpha)\mathcal{I}_{q}(\alpha)$. We now analyze each of these two functionals separately: using (4.11), (4.16), (4.18) and the fact that by the definition in (4.13),
$0\leq\mathcal{G}_{1}(\alpha,d\varphi)\leq 2\|d\varphi\|_{L^{1}({\mathbb{S}}^{3})}\leq 2\sqrt{2}\pi\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}\,,$
we find that
$\begin{split}\rm{(i)}\quad&\frac{\mathcal{I}_{1}(\alpha)}{Q(\alpha)^{\frac{1}{2}}}-\mathcal{I}_{1}(\alpha^{+}_{0,1})=-\frac{1}{4}\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi+\pi^{2}\big{(}Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]\big{)}^{2}\\ &\qquad\qquad\qquad\qquad+\bigg{(}1-\frac{1}{2}Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]\bigg{)}\mathcal{G}_{1}(\alpha^{+}_{0,1},d\varphi)+\mathcal{O}(\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{3})\,,\\[5.0pt] \rm{(ii)}\quad&\frac{\mathcal{I}_{2}(\alpha)}{Q(\alpha)}-\mathcal{I}_{2}(\alpha^{+}_{0,1})=\int_{\mathbb{S}^{3}}|d\varphi|^{2}-2\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi+\mathcal{O}(\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{3})\,.\end{split}$ (4.19)
Collecting (4.3), the expansions (4.19)(i)-(ii), and using also (4.17), we find precisely (4.3). ∎
@@
Let
$\alpha^{+}_{0,1}\in\underset{\widetilde{\alpha}\in E_{0,1}^{+}}{\operatorname{argmin}}\,\|\alpha-\widetilde{\alpha}\|_{L^{2}({\mathbb{S}}^{3})}\,.$ (4.20)
@@
$\alpha^{+}_{0,1}\in\underset{\widetilde{\alpha}\in E_{0,1}^{+}}{\operatorname{argmin}}\,\big{\|}P_{E_{0}^{+}}\alpha-\widetilde{\alpha}\big{\|}_{L^{2}({\mathbb{S}}^{3})}\,,$ (4.21)
where $P_{E_{0}^{+}}\alpha$ is the $L^{2}$-orthogonal projection of $\alpha$ onto $E_{0}^{+}$. Thus, $(P_{E_{0}^{+}}\alpha-\alpha^{+}_{0,1})$ is in the normal space to $E^{+}_{0,1}\cong\mathbb{S}_{4}^{2}\subset\mathbb{R}^{3}$ at $\alpha^{+}_{0,1}$, i.e.,
$P_{E_{0}^{+}}\alpha=(1+t)\alpha^{+}_{0,1}\,\ \text{for some }t\in\mathbb{R}\,.$ (4.22)
@@
$\frac{\alpha}{1+t}=\alpha^{+}_{0,1}+d\psi\,,\quad\psi:=\frac{\varphi-t\beta_{0}}{1+t}\,,$ (4.24)
where we note that, by (4.22), $d\psi\in(E_{0}^{+})^{\bot}$, the latter intended in the $L^{2}$-sense: hence
$\int_{{\mathbb{S}}^{3}}\langle\alpha^{+}_{0,1},d\psi\rangle=0\,.$ (4.25)
We have $\mathcal{E}_{\rho}(\frac{\alpha}{1+t})=\mathcal{E}_{\rho}(\alpha)$, and applying Lemma 4.3 to $\frac{\alpha}{1+t}$ and using (4.13) and (4.25), we find
$\begin{split}\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0,1})&=\rho^{-2}\bigg{(}\int_{\mathbb{S}^{3}}|d\psi|^{2}-2\int_{\mathbb{S}^{3}}\psi\wedge d\psi\bigg{)}-\frac{1}{2}\int_{\mathbb{S}^{3}}\psi\wedge d\psi\\ &\qquad+2\mathcal{G}_{1}(\alpha^{+}_{0,1},d\psi)+(2+\rho^{-2})\,\mathcal{O}\left(\|d\psi\|_{L^{2}({\mathbb{S}}^{3})}^{3}\right)\\ &\geq\rho^{-2}\bigg{(}\int_{\mathbb{S}^{3}}|d\psi|^{2}-2\int_{\mathbb{S}^{3}}\psi\wedge d\psi\bigg{)}-\frac{1}{2}\int_{\mathbb{S}^{3}}\psi\wedge d\psi\\ &\qquad+(2+\rho^{-2})\,\mathcal{O}\left(\|d\psi\|_{L^{2}({\mathbb{S}}^{3})}^{3}\right)\,.\end{split}$ (4.26)
We now proceed to give a lower bound on the last line of (4.26). Using Proposition 2.2 and that $d\psi\in(E_{0}^{+})^{\bot}$, since $H^{2}_{\rm{dR}}({\mathbb{S}}^{3})=0$ we can write
$d\psi=\sum_{k=0}^{\infty}(\eta_{k}^{+}+\eta_{k}^{-})\,,\ \text{where }\eta_{0}^{+}=0\,,\ \eta^{\pm}_{k}=d\psi^{\pm}_{k}\in E_{k}^{\pm}\,,$ (4.27)
where $\eta_{k}^{\pm}$ is the $L^{2}$-orthogonal projection of $d\psi$ onto $E_{k}^{\pm}$. Since for every $\ell\in\mathbb{N}$ the 2-form $d\psi^{\pm}_{\ell}$ solves (2.8) with eigenvalue $\lambda^{\pm}_{\ell}:=\pm(\ell+2)$, using also Lemma 2.1 we have that, for $k,\ell\in\mathbb{N}$,
In particular, (4.28) implies that for every $k\geq 1$,
$\rho^{-2}\bigg{(}\int_{\mathbb{S}^{3}}|d\psi_{0}^{-}|^{2}-2\int_{\mathbb{S}^{3}}\psi_{0}^{-}\wedge d\psi_{0}^{-}\bigg{)}-\frac{1}{2}\int_{\mathbb{S}^{3}}\psi_{0}^{-}\wedge d\psi_{0}^{-}$ $=\frac{8\rho^{-2}+1}{4}\int_{\mathbb{S}^{3}}|d\psi_{0}^{-}|^{2}\,,$
$\rho^{-2}\bigg{(}\int_{\mathbb{S}^{3}}|d\psi_{k}^{\pm}|^{2}-2\int_{\mathbb{S}^{3}}\psi_{k}^{\pm}\wedge d\psi_{k}^{\pm}\bigg{)}-\frac{1}{2}\int_{\mathbb{S}^{3}}\psi_{k}^{\pm}\wedge d\psi_{k}^{\pm}$ $=\bigg{(}\rho^{-2}\big{(}1-\frac{2}{\lambda_{k}^{\pm}}\big{)}-\frac{1}{2\lambda_{k}^{\pm}}\bigg{)}\int_{\mathbb{S}^{3}}|d\psi_{k}^{\pm}|^{2}$
$\geq\frac{2\rho^{-2}-1}{6}\int_{\mathbb{S}^{3}}|d\psi_{k}^{\pm}|^{2}\,,$ (4.29)
with equality holding in the last line for $\psi_{1}^{+}$. Hence, in view of the pairwise $L^{2}$-orthogonality of $(\eta_{k}^{\pm})_{k\in\mathbb{N}}$ and (4.28), we see that the quadratic form in the second line in (4.26) splits orthogonally with respect to the decomposition (4.27). Thus, using (4) we find
$\begin{split}\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0,1})&\geq\frac{2\rho^{-2}-1}{6}\int_{\mathbb{S}^{3}}|d\psi|^{2}+\left(2+\rho^{-2}\right)\mathcal{O}\left(\|d\psi\|_{L^{2}({\mathbb{S}}^{3})}^{3}\right)\\ &\geq\left(\frac{2\rho^{-2}-1}{6}-C\,\varepsilon_{0}\left(2+\rho^{-2}\right)\right)\int_{\mathbb{S}^{3}}|d\psi|^{2}\,.\end{split}$ (4.30)
To complete the proof it suffices to show that $\psi$ can be replaced by $\varphi$ in (4.30), as then we can choose
$0<\varepsilon_{0}\leq\frac{2\rho^{-2}-1}{12C(2+\rho^{-2})}\,.$
From the definition (4.24) of $\psi$, (4.10), (4.11) and (4.25), we have
$\int_{\mathbb{S}^{3}}|d\varphi|^{2}=(1+t)^{2}\int_{\mathbb{S}^{3}}|d\psi|^{2}+32\pi^{2}t^{2}\,,$ (4.31)
and as we show next
$|t|=\mathcal{O}\big{(}\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{2}\big{)}\,.$ (4.32)
Indeed, since $Q(\alpha)=Q(\alpha^{+}_{0,1})=1$, by (2.5) we have
@@
where in the last equality we used again (4.10), (4.11), (4.24), and (4.25). Thus, $t=-\frac{1}{2}Q(d\varphi)$ and so using also (2.6), (4.32) follows. The latter, together with the assumption (4.7), which can be rewritten as $\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}\leq\varepsilon_{0}$, imply that we can assume $|t|\leq\tfrac{1}{2}$, provided that $\varepsilon_{0}>0$ is chosen sufficiently small, which already shows that $\mathcal{O}(\|d\psi\|^{3}_{L^{2}({\mathbb{S}}^{3})})$ can be replaced by $\mathcal{O}(\|d\varphi\|^{3}_{L^{2}({\mathbb{S}}^{3})})$ in (4.30). Finally, again by (4.31) and (4.32), we can estimate
$\int_{\mathbb{S}^{3}}|d\psi|^{2}\geq\frac{\int_{\mathbb{S}^{3}}|d\varphi|^{2}-\tfrac{C^{2}}{32\pi^{2}}\big{(}\int_{{\mathbb{S}}^{3}}|d\varphi|^{2})^{2}}{(1+t)^{2}}\geq\frac{1}{3}\int_{{\mathbb{S}}^{3}}|d\varphi|^{2}\,,$
where $C>0$ is the constant of (2.6), and $\varepsilon_{0}$ is small enough so that $0<\varepsilon_{0}\leq\frac{2\sqrt{2}\pi}{C}$. Hence (4.30) also holds with $d\psi$ replaced by $d\varphi$, which completes the proof of the proposition. ∎
5. Proof of Theorem 1.1
In [31] it is shown that elements in $E^{+}_{0,1}$ minimize the functional $\alpha\mapsto\mathcal{I}_{2}(\alpha)/Q(\alpha)$ on $\Omega^{2}(\mathbb{S}^{3})$. Actually, we can obtain an associated quantitative estimate, as follows.
Proposition 5.1.
$\mathcal{I}_{2}(\alpha)-32\pi^{2}Q(\alpha)\geq\frac{1}{4}\min_{\widetilde{\alpha}^{+}_{0}\in E_{0}^{+}}\int_{\mathbb{S}^{3}}|\alpha-\widetilde{\alpha}^{+}_{0}|^{2}\,.$ (5.1)

### v2

@@ 4. Relaxed form of the perturbed Faddeev–Skyrme energy
In [speight2011] it is shown that elements in $E^{+}_{0,1}$ minimize the functional $\alpha\mapsto\mathcal{I}_{2}(\alpha)/Q(\alpha)$ on $\Omega^{2}(\mathbb{S}^{3})$. Actually, we can obtain an associated quantitative estimate, for both the eigenspaces $E_{0}^{+}$ and $E_{0}^{-}$, as follows.
Proposition 4.3.
The following estimates hold true:
  • $\rm{(i)}$
    For every $\psi\in W^{1,2}({\mathbb{S}}^{3};\bigwedge^{1}T^{*}{\mathbb{S}}^{3})$ such that $d\psi\bot_{L^{2}}\,E_{0}^{+}\oplus E_{0}^{-}$, it holds that
    $\Big|\int_{{\mathbb{S}}^{3}}\psi\wedge d\psi\Big|\leq\frac{1}{3}\int_{{\mathbb{S}}^{3}}|d\psi|^{2}\,.$ (4.10)
  • $\rm{(ii)}$
    $\mathcal{I}_{2}(\alpha)\mp 32\pi^{2}Q(\alpha)\geq\frac{1}{3}\min_{\widetilde{\alpha}^{\pm}_{0}\in E_{0}^{\pm}}\int_{\mathbb{S}^{3}}|\alpha-\widetilde{\alpha}^{\pm}_{0}|^{2}\,.$
Proof.
For (i), using Proposition 2.2, the $L^{2}$-orthogonality assumption, and since $H^{2}_{\rm{dR}}({\mathbb{S}}^{3})=0$, we can write
$d\psi=\sum_{k=0}^{\infty}(\eta_{k}^{+}+\eta_{k}^{-})\,,\ \text{where }\eta_{0}^{\pm}=0\,,\ \eta^{\pm}_{k}=d\psi^{\pm}_{k}\in E_{k}^{\pm}\,$,
where $\eta_{k}^{\pm}$ is the $L^{2}$-orthogonal projection of $d\psi$ onto $E_{k}^{\pm}$. Since for every $\ell\in\mathbb{N}$ the 2-form $d\psi^{\pm}_{\ell}$ solves (2.8) with eigenvalue $\lambda^{\pm}_{\ell}:=\pm(\ell+2)$, using also Lemma 2.1(i), we have that for $k,\ell\in\mathbb{N}$,
In particular, (4.11) together with the fact that $|\lambda^{\pm}_{k}|\geq 3$ for $k\geq 1$, indeed implies
$\Big|\int_{{\mathbb{S}}^{3}}\psi\wedge d\psi\Big|$ $=\Big|\sum_{k,\ell\geq 1}\int_{{\mathbb{S}}^{3}}\psi_{k}^{\pm}\wedge d\psi_{\ell}^{\pm}\Big|\leq\sum_{k\geq 1}\frac{1}{|\lambda^{\pm}_{k}|}\int_{{\mathbb{S}}^{3}}|d\psi_{k}^{\pm}|^{2}\leq\frac{1}{3}\int_{{\mathbb{S}}^{3}}|d\psi|^{2}\,$,
with equality holding for $\psi_{1}^{\pm}$.
For the proof of (ii), consider a closed $2$-form $\alpha$, thus $\alpha=d\beta$ for some $\beta\in W^{1,2}({\mathbb{S}}^{3};\bigwedge^{1}T^{*}{\mathbb{S}}^{3})$ with $d^{*}\beta=0$. The lowest eigenvalue of the Hodge Laplacian $\Delta:=dd^{*}+d^{\ast}d$ on closed $2$-forms is $4$, and the corresponding eigenspace is given by
$\ker d\cap\ker(\Delta-4)=E_{0}^{+}\oplus E_{0}^{-}\,$,
cf. [isobe2008, Section 3] and (2.9). Therefore, we can ($L^{2}$-orthogonally) decompose
$\beta=\alpha_{0}^{+}+\alpha_{0}^{-}+d\varphi\,,\ \text{where }\ d*\alpha_{0}^{\pm}=\pm 2\alpha_{0}^{\pm}\,,\ \ d\varphi\perp_{L^{2}}E_{0}^{+}\oplus E_{0}^{-}\,$, (4.12)
Using (4.12) and (4.13), we find
$\mathcal{I}_{2}(\alpha)=4\int_{\mathbb{S}^{3}}|\alpha_{0}^{+}|^{2}+4\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\int_{\mathbb{S}^{3}}|d\psi|^{2}\,$, (4.14)
and
$Q(\alpha)=\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\langle*\beta,\alpha\rangle\\ =\frac{1}{8\pi^{2}}\int_{\mathbb{S}^{3}}|\alpha_{0}^{+}|^{2}-\frac{1}{8\pi^{2}}\int_{\mathbb{S}^{3}}|\alpha_{0}^{-}|^{2}+\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\psi\wedge d\psi\,.$ (4.15)
We now proceed to give a lower bound on the last term in the right hand side of (4.15). Combining (4.14) and (4.15), using the estimate (4.10) and again (4.13), we deduce
[…]
concluding the proof. ∎
The next lemma gives us a precise expansion for the energy $\mathcal{E}_{\rho}$ around elements in $E^{+}_{0,1}$.
@@
[…]
where we have set
@@
$\mathcal{E}_{\rho}(\alpha)=\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\mathcal{I}_{1}(\alpha)+\rho^{-2}\mathcal{I}_{2}(\alpha)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,.$ (4.18)
We first expand the two quantities appearing in (4.18) separately, obtaining:
[…]
@@
$\mathcal{I}^{\prime}_{q}(\alpha^{+}_{0,1})[d\varphi]:=\frac{d}{dt}\Big|_{t=0}\mathcal{I}_{q}(\alpha^{+}_{0,1}+td\varphi)\,$,
and similarly for $Q^{\prime}(\alpha^{+}_{0,1})[d\varphi]$, which we will also compute here for later purposes. Note that (4.19)(i) follows simply by the definition of $\mathcal{G}_{1}$ in (4.17), and (4.19)(ii) by a simple expansion of the quadratic energy $\mathcal{I}_{2}$.
To proceed further we note that, by the definition of $\beta_{0}$ in (4.8) and Lemma 2.1(i), we have
@@
[…]
this last equation, together with $Q(\alpha^{+}_{0,1})=1$, expresses the fact that $\alpha^{+}_{0,1}$ is a critical point of the functionals $\alpha\mapsto Q^{-q/2}(\alpha)\mathcal{I}_{q}(\alpha)$. We now analyze each of these two functionals separately: using (4.9) and (4.19)-(4.21), we find that
$\begin{split}\rm{(i)}\quad&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}_{1}(\alpha)-\mathcal{I}_{1}(\alpha^{+}_{0,1})=\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{4}\int_{{\mathbb{S}}^{3}}\langle\alpha^{+}_{0,1},d\varphi\rangle+\mathcal{G}_{1}(\alpha^{+}_{0,1},d\varphi)\,,\\[5.0pt] \rm{(ii)}\quad&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}_{2}(\alpha)-\mathcal{I}_{2}(\alpha^{+}_{0,1})=\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{{\mathbb{S}}^{3}}|d\varphi|^{2}+2\int_{{\mathbb{S}}^{3}}\langle\alpha^{+}_{0,1},d\varphi\rangle\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,.\end{split}$ (4.22)
Collecting (4.2), and the expansions (4.22)(i)-(ii), we find precisely (4.16). ∎
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha^{+}_{0,1}\in\underset{\widetilde{\alpha}\in E_{0,1}^{+}}{\mathrm{argmin}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\big\|\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{0}^{+}-\widetilde{\alpha}\big\|_{L^{2}({\mathbb{S}}^{3})}\,$,
[…]
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha_{0}^{+}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=(1+t)\alpha^{+}_{0,1}\,\ \text{for some }t\in\mathbb{R}\,.$
@@
[…]
where we note that $d\psi\in(E_{0}^{+})^{\bot}$, the latter intended in the $L^{2}$-sense: hence, by (4.24) and using also (4.9), we have
$\int_{{\mathbb{S}}^{3}}\langle\alpha^{+}_{0,1},d\psi\rangle=0\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\iff t=\frac{1}{32\pi^{2}}\int_{{\mathbb{S}}^{3}}\langle\alpha^{+}_{0,1},d\varphi\rangle\,.$ (4.25)
Moreover, since $Q(\alpha)=Q(\alpha^{+}_{0,1})=1$, by (2.5) we have
@@
where in the last equality we used again (4.8), (4.9), and (4.25). Thus,
$\frac{1}{32\pi^{2}}\int_{{\mathbb{S}}^{3}}\langle\alpha^{+}_{0,1},d\varphi\rangle=t=-\frac{1}{2}Q(d\varphi)\,.$ (4.26)
Note further that by (4.5), and since $\alpha^{+}_{0,1}\in E_{0,1}^{+}$, we have
$\|\alpha-\alpha^{+}_{0,1}\|_{L^{2}({\mathbb{S}}^{3})}\leq\|\alpha+\alpha^{+}_{0,1}\|_{L^{2}({\mathbb{S}}^{3})}\iff\int_{{\mathbb{S}}^{3}}\langle\alpha,\alpha^{+}_{0,1}\rangle\geq 0\,$,
which, by (4.23) is equivalent to the simple bound
$t\geq-1\iff Q(d\varphi)\leq 2\,.$ (4.27)
The definition of $\psi$ in (4.24), together with (4.25)-(4.27) further yield the relation
[…]
since (4.27) and the first line in the above identity imply that $Q(d\psi)\leq 1$.
We can now apply Lemma 4.4 and use (4.17), (4.24), and (4.26), to find
$\begin{split}\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0,1})&\geq\rho^{-2}\int_{\mathbb{S}^{3}}|d\varphi|^{2}-16\pi^{2}\Big(\frac{1}{2}+\frac{2}{\rho^{2}}\Big)Q(d\varphi)\\ &=\rho^{-2}\Big(\mathcal{I}_{2}(d\varphi)-32\pi^{2}Q(d\varphi)\Big)-8\pi^{2}Q(d\varphi)\\ &=\rho^{-2}\Big(\mathcal{I}_{2}(d\psi)-32\pi^{2}Q(d\psi)\Big)-16\pi^{2}\big(1-\sqrt{1-Q(d\psi)}\big)\,.\end{split}$ (4.29)
The last equality follows from (4.28), the fact that
$\mathcal{I}_{2}(d\varphi)=\mathcal{I}_{2}(d\psi)+t^{2}\mathcal{I}_{2}(\alpha_{0,1}^{+})$
by the orthogonality in (4.25), and that
$\mathcal{I}_{2}(\alpha_{0,1}^{+})=32\pi^{2}Q(\alpha^{+}_{0,1})=32\pi^{2}\,$,
cf. (4.9). We rewrite (4.29) as
[…]
$\quad+\frac{1}{2}\Big(\mathcal{I}_{2}(d\psi)-48\pi^{2}Q(d\psi)\Big)\,.$
Note that in view of (4.1)-(4.10) combined with decomposing $d\psi\in(E_{0}^{+})^{\perp}$ according to the orthogonal sum $(E_{0}^{+})^{\perp}=(E_{0}^{+}\oplus E_{0}^{-})^{\perp}\oplus E_{0}^{-}$, the last line in the above inequality is nonnegative, hence we deduce
[…]
$\quad-16\pi^{2}\Big(1-\sqrt{1-Q(d\psi)}-\frac{Q(d\psi)}{2}\Big)\,$, (4.30)
which can be equivalently written as,
$\begin{split}\frac{\mathcal{E}_{\rho}(\alpha)-\mathcal{E}_{\rho}(\alpha^{+}_{0,1})}{16\pi^{2}}&\geq(2\rho^{-2}-1)\Big(\frac{\mathcal{I}_{2}(d\psi)}{32\pi^{2}}-Q(d\psi)\Big)-\frac{1}{2}|Q(d\psi)|\\ &\quad+\frac{1}{2}|Q(d\psi)|-\Big(1-\sqrt{1-Q(d\psi)}-\frac{Q(d\psi)}{2}\Big)\,.\end{split}$ (4.31)
By the elementary fact that the smallest value of $\lambda>0$ such that the function
$(-\infty,1]\ni s\mapsto\lambda|s|-\Big(1-\sqrt{1-s}-\frac{s}{2}\Big)\,$,
is nonnegative is $\lambda=1/2$, (4.31) together with (4.1) and (4.10) again, finally yield,
[…]
the latter inequality holding for $0<\rho\leq 1$. This proves the desired minimality (4.6) of $\alpha_{0,1}^{+}$.
[…]
Remark 4.5 (On the threshold for linear stability).
A similar spectral analysis as the one in the proof of Proposition 4.3 had already been used by Isobe in [isobe2008] to prove that
[…]
and by continuity weakly stable also for $\rho_{\ast}=\sqrt{2}$, cf. the first assertion of Theorem 1.1, Proposition 4.1 and Corollary 4.1 therein. The role of the threshold $\rho_{*}=\sqrt{2}$ for the linear stability of $h$ is also revealed by our calculations in the proof Proposition 4.2, for instance in (4.30). Indeed, by Proposition 4.3 and that
$1+\theta\geq 2\sqrt{\theta}\ \quad\forall\theta\geq 0\,$,
the latter applied for $\theta:=1-Q(d\psi)$, if $\rho>\sqrt{2}$ the right hand side of (4.30) becomes negative, leading to an insufficient lower bound for the conclusion.
As an immediate consequence of Proposition 4.2, we obtain the minimality of the Hopf map for the $\mathcal{FS}_{\rho}$ in the claimed range.
Corollary 4.6.
For every $\rho\in(0,1]$ and $u\in\mathcal{U}_{\mathcal{FS}}$ with $Q(u)=1$, it holds that
$\mathcal{FS}_{\rho}(u)\geq\mathcal{FS}_{\rho}(h)\,$,
with equality if and only if $u^{*}\omega_{\mathbb{S}^{2}}\in E_{0,1}^{+}$.